OpenStudy (anonymous):

The indefinite integral of [1]/[x^2 + 4] dx Attempt: Is this not arctan(x/2)*2?

OpenStudy (zarkon):

divided by 2 not times 2

OpenStudy (anonymous):

Why?

OpenStudy (zarkon):

\[\frac{d}{dx}arctan(x/2)=\frac{1}{(x/2)^2+1}\frac{1}{2}\] \[=\frac{1}{x^2/4+1}\frac{1}{2}=\frac{1}{x^2/2+2}\] if we multiply by 1/2 we get the original integrand

OpenStudy (anonymous):

I understand. Thank you.